OpenStudy (anonymous):

What are the solutions of the equation (x + 2)2 – 2(x + 2) – 15 = 0? Use u substitution to solve. (Will get medal)

OpenStudy (anonymous):

set u=(x+2) \[u ^{2}-2u-15=0\] let's concentrate on just the left side We can solve this by factoring: x² - 2x - 15 = 0 x² - 5x + 3x - 15 = 0 x(x - 5) + 3(x - 5) = 0 (x - 5)(x + 3) = 0 x - 5 = 0 → x = 5 x + 3 = 0 → x = -3 We can also solve this, by completing the square: x² - 2x - 15 = 0 x² - 2x = 15 x² - 2x + (-1)² = 15 + (-1)² x² - 2x + 1 = 15 + 1 x² - 2x + 1 = 16 (x - 1)(x - 1) = 16 √(x - 1)² = ±16 x - 1 = ±4 x - 1 = 4 → x = 5 x - 1 = -4 → x = -3 did the rest in x in case you didn't notice but u=2 and u=-3 and for the last part u=x+2 so to find the real x x=2-u plug in your u and you'll find your two answers.